The length of a rectanglar garden is 4 m greater than the width. The area of the garden is 96 cm.sq. Find the length and width.
Answer by checkley77(12844)

L=W+4
A=LW
96=(W+4)W
96=W^2+4W
W^2+4W-96=0
(W+12)(W-8)=0
W-8=0
W=8 ANS.
L=8+4=12 ANS.
PROOF:
96=12*8
96=96